# The NANOGrav 15 year Data Set: Removing pulsars one by one from the pulsar timing array

### v1

@@ 4.1 The full array
We begin by calculating the signal found in the timing residuals of the full NG15 timing array comprised of 67 pulsars.
The results from both the single-component (SC) NMOS and the multiple-component (MC) NMOS analyses are shown in Table 1 and Figure 1, and were calculated for monopole, dipole and HD spatial correlations. Both our SCNMOS and MCNMOS results for the full array are consistent with those of NG15 (Figure 4 and Table 2 in Agazie et al. (2023a) respectively). As we are using the full array in this case, this also permits us to compare our results with the fully Bayesian analysis which includes HD spatial correlations in the Agazie et al. (2023a) paper. Again they are in agreement.
@@
Table 1: Mean values for the optimal statistic $\hat{A}$ and S/N for the three ORFs and for both SCNMOS and MCNMOS, taken from Figure 1. These are calculated when all 67 NANOGrav pulsars are included in the pulsar timing array. The mean (and not the median) is used here to allow direct comparison with values calculated in NG15.
@@
Figure 1: Top plot: Results for the single-component NMOS (SCNMOS). Bottom plot: Results for the multiple-component NMOS (MCNMOS). In both cases the narrow top panels shows the distribution of the squared amplitudes $\hat{A}^{2}$ for the monopole, dipole and HD components while the wider bottom panels show the distribution of the S/N for the same three spatial correlations. The values were calculated 10,000 times as part of the noise marginalization process. Dashed vertical lines indicate the mean value for a distribution. The mean (and not the median) is used here to allow direct comparison with values calculated in NG15.
4.2 Removing pulsars: single-component analysis
@@
We next remove the least-noisy pulsars from the full PTA one by one. As seen in the upper panel of Figure 2 the SCNMOS for the dipole and monopole spatial correlation is around $10^{-30}$ and remains relatively stable until approximately half of the pulsars have been excised from the full PTA. Upon the removal of the 37th pulsar (PSR J2317$+$1439) the dipole spatial correlation OS suddenly increases. The SCNMOS for the HD spatial correlation remains relatively flat until approximately a third of the pulsars have been removed. At this point it experiences a steady decline until an abrupt increase upon the removal of the 34th pulsar (PSR B1855$+$09), followed by an equally abrupt drop as the next pulsar is removed. In the lower panel of Figure 2 all three spatial correlations begin with a S/N of around 4 and then undergo a relatively steady drop to a S/N around zero after the removal of 33 pulsars. However, there is a significant drop in S/N (especially for HD and monopole spatial correlations) when the first pulsar is removed. The jumps seen in the optimal statistic after the removal of pulsar 34 and 37 are reflected in the S/N for the HD and dipole spatial correlations respectively.
@@
Figure 2: The evolution of the SCNMOS squared amplitude $\hat{A}^{2}$ and its S/N for NG15 as the least-noisy pulsars are removed one by one from the PTA. The bottom axis is the number of pulsars removed while the top axis gives the name of most recent pulsar to have been removed. Each circle in the upper panel represents the median of the 10,000 SCNMOS $\hat{A}^{2}$ distribution calculated for a specific subset of the entire PTA. The bars attached to the circles represent the range between the 5th and 95th percentiles of the $\hat{A}^{2}$ distribution. In the lower panel the circles and bars depict the median and range respectively of the S/N values of the SCNMOS analysis. Note that one cannot assess the significance in changes in the median circles by referring to the bars depicting the spread of the distribution from which it was calculated; the uncertainty on the median is of order 100 ($\sqrt{10,000}$) times smaller than the width of the range bars. Data within the gray-shaded region are not considered when calculating the metrics defined in Section 5.1 as the behavior of $\hat{A}^{2}$ and S/N can become increasingly erratic due the small number of low-sensitivity pulsars remaining.
4.3 Removing pulsars: multiple-component analysis
@@
As expected, the main features described above and seen in Figure 2 also appear in Figure 3. The S/N values, however, are systematically lower for the multiple-component analysis compared to the single-component analysis.
@@
Figure 3: The evolution of the MCNMOS and S/N for NG15 as the least-noisy pulsars are removed one by one from the PTA. Similar to Figure 2; see that caption for more details.
@@
As the MCOS more accurately recovers spatially correlated signals than the OS (Sardesai et al., 2023), we have identified four features of interest in the MCNMOS pulsar-removal plots of Figure 3. The features are described and quantified in the following and are highlighted in Figure 4. Although we focused on these four features, it is important to note that various others could have been selected; we primarily chose these particular features based on notable characteristics observed in the plotted data.
@@
Feature (I): Initial S/N for the HD spatial correlations. The median S/N of the signal with HD spatial correlations when all 67 NANOGrav pulsars are included in the pulsar timing array is $\sim$2.8. We see a value close to this in Table 2 of Agazie et al. (2023a).
@@
Feature (II): S/N drop after removal of one pulsar for the HD spatial correlations. When we remove the least-noisy pulsars first, a S/N drop from $\sim$2.8 to $\sim$2.0 is seen in the HD spatial correlations plot when the first pulsar PSR J1909--3744 is removed. This drop of 29% after the removal of only one pulsar from the PTA appears notably large. In the highly idealized case where all pulsars are identical, the S/N is expected to scale linearly with the number of pulsars in the array  (see e.g., Equation 24 from Siemens et al., 2013); a 29% drop is clearly much larger than this naive approximation would predict. However, PSR J1909--3744 is the most precisely timed pulsar and so its removal would be expected to produce a larger drop in S/N compared to other pulsars. Specifically, the weighted root-mean-square of epoch-averaged post-fit timing residuals after whitening for PSR J1909--3744 is 0.066 $\mu$s (Agazie et al., 2023b), which is $\sim$43 times smaller than the highest value in NG15. Nevertheless, the magnitude of the drop still seems surprising for a single pulsar removal and its significance will be quantified later in this paper.
@@
Feature (III): Sharp Increase in S/N for HD spatial correlations. When 33 pulsars have been removed in the order of increasing noise, the S/N of the HD spatial correlation is $\sim$-0.1. On the removal of the 34th pulsar (PSR B1855$+$09) the S/N for the PTA consisting of the remaining 33 pulsars rises to $\sim$1.4. A priori the S/N would be expected to decrease when a pulsar is removed. It is surprising, therefore, that the removal of a single pulsar could produce such a significant S/N increase.
@@
Feature (IV): Sharp Increase in S/N for dipole spatial correlations. The S/N of the dipole spatial correlation jumps from $\sim$1 to $\sim$3 when the 37th pulsar (PSR J2317+1439) is removed in the order of increasing noise.
@@ 5 Comparison with simulated data
To assess the peculiarity of the prominent features we see in our MCNMOS analysis of NG15, we make a comparison with an ensemble of simulated data sets that share the statistical properties of NG15. The simulated data were generated in Agazie et al. (2023a) following Pol et al. (2021). We note that the construction of these simulated data sets is central to the conclusions of this paper. Comprehensive details on their construction can be found in these previous references, but to summarize, each of the 100 simulated data sets share the same pulsars, timing cadence and noise parameters as NG15 and also include a spatially correlated power-law SGWB with power-law amplitude $2.7\times 10^{-15}$ and spectral index $-13/3$. The data sets differ only in their random noise realizations. For each realization, we calculated the MCNMOS and removed pulsars one by one as in the analysis of NG15. By performing the same analysis on simulated data sets we are able to determine the expected behavior of the MCNMOS as pulsars are steadily removed. By comparing the results from the NG15 data with those from the ensemble of simulated data it is possible to assign a significance to the unexpected features identified in Section 4.3. Fig. 6 shows the results for a single representative simulated data set with the least-noisy pulsars removed first; this can be compared to the equivalent plot from the NG15 data set, shown in Fig. 3.
@@ 5.1 Defining metrics
The metrics focus on S/N because each of our four features in Section 4.3 appears in S/N analyses. For completeness, however, we compute these metrics for $\hat{A}$ as well as the S/N. For each of these two variables, we calculate the three metrics for each of the three ORFs, resulting in a total of 18 metric values. These 18 values are calculated for NG15 and for the 100 simulations. This allows us to compute $p$-values thereby determining the probability that a given NG15 pulsar-removal feature would appear, given the null hypothesis that NG15 is accurately modeled and internally consistent. The 18 $p$-values are shown in Table 2; 3 of the 18 are less than or equal to 0.01. When calculating the metrics we only consider the data until the point where 50 pulsars have been removed; this is indicated in Figures 2, 3 and 6 by the gray regions. This cutoff is somewhat arbitrary, but is justified by the fact that after this point, the behavior of $\hat{A}^{2}$ and S/N becomes increasingly erratic because of the small number of low-sensitivity pulsars remaining. This is judged by eye from e.g., Figures 2, 3 and 4.
@@
Figure 4: The top panel shows the length of the timing baseline for each NG15 pulsar. The lower six panels are in pairs for each spatial correlation, tracking the MCNMOS and the S/N as the least-noisy pulsars are removed one by one from the PTA. The panels with a red line correspond to monopole spatial correlations. The panels with a green line correspond to dipole spatial correlations. The panels with a blue line correspond to HD spatial correlations. The coloured lines are the results from NG15 and each grey trace corresponds to one of the 100 simulated data sets. The black dashed lines shows the median of the 100 grey traces. The features described in Section 4.3 are highlighted.
@@ 6 Discussion
The behavior of the MCNMOS as applied to NG15, shown in Figure 3 and described in Section 4, exhibits several initially unexpected features as pulsars are steadily removed from the array. To contextualize the features and to determine whether they point to some unmodeled processes in the NG15, we have repeated the analysis using simulated data with known embedded GW signal and noise properties. To quantify the significance of the unexpected NG15 features we calculate $p$-values for 18 applications of the three metrics described in Section 5.1. Out of 18 instances, there are 3 with with $p$-values less than or equal to 0.01.
@@
Dropout analysis (Aggarwal et al., 2019; Arzoumanian et al., 2020) was employed in NG15 to evaluate the significance of the GWB signal in individual pulsars. The dropout method introduces a binary parameter to either include or exclude the common signal for a specific pulsar. The dropout factor or the Bayes factor then indicates the degree to which each pulsar supports the presence of the GWB signal. In our approach, instead of evaluating each pulsar independently we removed pulsars in a cumulative manner. While both methods aim to understand the robustness and consistency of the detected signal, the dropout method focuses on the individual contribution of each pulsar in isolation. In contrast, our cumulative removal approach provides insight into the collective impact on the signal as pulsars are removed and so no single removal (other than the first) can be viewed in isolation. Comparing Figure 2 with Figure 7 and Figure 3 with Figure 8 we see that the effect of removing any particular pulsar can be quite different depending on the order of removal. Furthermore, Figure 4 demonstrates that even when the order of removal is preserved, the removal of a particular pulsar can have a different effect on the recovered signal for different realizations of the simulation.
@@ 7 Conclusions and Future Work
We have investigated the behavior of recovered signal strength when pulsars are removed one by one from the NANOGrav 15-year data set. Despite variations during pulsar removal that we had not anticipated, these fluctuations are consistent with those observed in simulated data sets that have matching noise properties and a gravitational wave background injected. After calculating $p$-values for the data we find that the results are consistent with the simulated data sets and, therefore, with a stochastic gravitational wave background signal. The purpose of our analysis was to provide internal consistency checks on the NANOGrav 15-year data set, and the results indicate that the data have successfully passed these tests. It also remains essential to conduct further investigations that may reveal systematics or unmodeled physical effects within the data.
The tools developed here have proved to be useful for checking the internal consistency of PTA data. Although we have applied them to NANOGrav 15-year data set in this work, these tools could be applied routinely to new PTA data sets as they become available. We plan to analyse the EPTA2 data set next, with an eye on IPTA DR3 and MeerKAT data.
@@
An alphabetical-order author list was used for this paper in recognition of the fact that a large, decade-timescale project such as NANOGrav is necessarily the result of the work of many people. All authors contributed to the activities of the NANOGrav Collaboration leading to the work presented here and reviewed the manuscript, text, and figures prior to the paper’s submission. Additional specific contributions to this paper are as follows. P.R.B., H.M., C.J.M., and A.V. collaboratively developed the concept and main ideas for this work. P.R.B. was responsible for data analysis and creating the figures. P.R.B., J.S.H., H.M., C.J.M., S.R.T., A.V., S.J.V., and C.A.W. contributed to writing the manuscript. N.S.P. created the simulated PTA data sets. G.A., A.A., A.M.A., Z.A., P.T.B., P.R.B., H.T.C., K.C., M.E.D., P.B.D., T.D., E.C.F., W.F., E.F., G.E.F., N.G., P.A.G., J.G., D.C.G., J.S.H., R.J.J., M.L.J., D.L.K., M. K., M.T.L., D.R.L., J.L., R.S.L., A.M., M.A.M., N.M., B.W. M., C.N., D.J.N., T.T.P., B.B.P.P., N.S.P., H.A.R., S.M.R., P. S.R., A.S., C.S., B.J.S., I.H.S., K.S., A.S., J.K.S., and H.M.W. developed the 15 yr data set through a combination of observations, arrival time calculations, data checks and refinements, and timing model development and analysis; additional specific contributions to the data set are summarized in Agazie et al. (2023b). J.S.H., J.G., N.L., N.S.P., J.P.S., and J.K.S. performed noise analyses on the data set.

### v2

@@ 4.1 The Full Array
We begin by calculating the signal found in the timing residuals of the full NG15 timing array comprised of 67 pulsars. The results from both the single-component SCNMOS and the MCNMOS analyses are shown in Table 1 and Figure 1, and were calculated for monopole, dipole and HD spatial correlations. Both our SCNMOS and MCNMOS results for the full array are consistent with those of NG15 (Figure 4 and Table 2 in Agazie et al. (2023a) respectively). As we are using the full array in this case, this also permits us to compare our results with the fully Bayesian analysis that includes HD spatial correlations in the Agazie et al. (2023a) paper. Again they are in agreement.
@@
Table 1: Mean values for the OS $\hat{A}$ and S/N for the three ORFs and for both SCNMOS and MCNMOS, taken from Figure 1. These are calculated when all 67 NANOGrav pulsars are included in the PTA. The mean (and not the median) is used here to allow direct comparison with the values calculated in NG15.
@@
Figure 1: Top plot: results for the SCNMOS. Bottom plot: results for the MCNMOS. In both cases the narrow top panels shows the distribution of the squared amplitudes $\hat{A}^{2}$ for the monopole, dipole and HD components, while the wider bottom panels show the distribution of the S/N for the same three spatial correlations. The values were calculated 10,000 times as part of the noise marginalization process. Dashed vertical lines indicate the mean value for a distribution. The mean (and not the median) is used here to allow direct comparison with values calculated in NG15.
4.2 Removing pulsars: SC analysis
@@
We next remove the least noisy pulsars from the full PTA one by one. As seen in the upper panel of Figure 2 the SCNMOS for the dipole and monopole spatial correlation is around $10^{-30}$ and remains relatively stable until approximately half of the pulsars have been excised from the full PTA. Upon the removal of the 37th pulsar (PSR J2317$+$1439) the dipole spatial correlation OS suddenly increases. The SCNMOS for the HD spatial correlation remains relatively flat until approximately a third of the pulsars have been removed. At this point it experiences a steady decline until an abrupt increase upon the removal of the 34th pulsar (PSR B1855$+$09), followed by an equally abrupt drop as the next pulsar is removed. In the lower panel of Figure 2 all three spatial correlations begin with an S/N of around 4 and then undergo a relatively steady drop to an S/N around zero after the removal of 33 pulsars. However, there is a significant drop in S/N (especially for HD and monopole spatial correlations) when the first pulsar is removed. The jumps seen in the OS after the removal of pulsar 34 and 37 are reflected in the S/N for the HD and dipole spatial correlations respectively.
@@
Figure 2: The evolution of the SCNMOS squared amplitude $\hat{A}^{2}$ and its S/N for NG15 as the least noisy pulsars are removed one by one from the PTA. The bottom axis is the number of pulsars removed while the top axis gives the name of most recent pulsar to have been removed. Each circle in the upper panel represents the median of the 10,000 SCNMOS $\hat{A}^{2}$ distribution calculated for a specific subset of the entire PTA. The bars attached to the circles represent the range between the 5th and 95th percentiles of the $\hat{A}^{2}$ distribution. In the lower panel the circles and bars depict the median and range respectively of the S/N values of the SCNMOS analysis. Note that one cannot assess the significance of the changes in the median circles by referring to the bars depicting the spread of the distribution from which it was calculated; the uncertainty on the median is of order 100 ($\sqrt{10,000}$) times smaller than the width of the range bars. The data within the gray-shaded region are not considered when calculating the metrics defined in Section 5.1 as the behavior of $\hat{A}^{2}$ and S/N can become increasingly erratic due the small number of low-sensitivity pulsars remaining.
4.3 Removing pulsars: MC analysis
@@
As expected, the main features described above and seen in Figure 2 also appear in Figure 3. The S/N values, however, are systematically lower for the MC analysis compared to the SC analysis.
@@
Figure 3: The evolution of the MCNMOS and S/N for NG15 as the least noisy pulsars are removed one by one from the PTA. Similar to Figure 2; see that caption for more details.
@@
As the MCOS more accurately recovers spatially correlated signals than the OS (Sardesai et al., 2023), we have identified four features of interest in the MCNMOS pulsar removal plots of Figure 3. The features are described and quantified in the following and are highlighted in Figure 4. Although we focused on these four features, it is important to note that various others could have been selected; we primarily chose these particular features based on notable characteristics observed in the plotted data.
@@
Feature (I). Initial S/N for the HD spatial correlations. The median S/N of the signal with HD spatial correlations when all 67 NANOGrav pulsars are included in the pulsar timing array is $\sim$2.8. We see a value close to this in Table 2 of Agazie et al. (2023a).
@@
Feature (II). S/N drop after removal of one pulsar for the HD spatial correlations. When we remove the least noisy pulsars first, an S/N drop from $\sim$2.8 to $\sim$2.0 is seen in the HD spatial correlations plot when the first pulsar PSR J1909--3744 is removed. This drop of 29% after the removal of only one pulsar from the PTA appears notably large. In the highly idealized case where all pulsars are identical, the S/N is expected to scale linearly with the number of pulsars in the array  (see e.g., Equation 24 from Siemens et al., 2013); a 29% drop is clearly much larger than this naive approximation would predict. However, PSR J1909--3744 is the most precisely timed pulsar and so its removal would be expected to produce a larger drop in S/N compared to other pulsars. Specifically, the weighted root-mean-square of epoch-averaged post-fit timing residuals after whitening for PSR J1909--3744 is 0.066 $\mu$s (Agazie et al., 2023b), which is $\sim$43 times smaller than the highest value in NG15. Nevertheless, the magnitude of the drop still seems surprising for a single pulsar removal and its significance will be quantified later in this paper.
@@
Feature (III). Sharp Increase in S/N for HD spatial correlations. When 33 pulsars have been removed in the order of increasing noise, the S/N of the HD spatial correlation is $\sim$-0.1. On the removal of the 34th pulsar (PSR B1855$+$09) the S/N for the PTA consisting of the remaining 33 pulsars rises to $\sim$1.4. A priori the S/N would be expected to decrease when a pulsar is removed. It is surprising, therefore, that the removal of a single pulsar could produce such a significant S/N increase.
@@
Feature (IV). Sharp Increase in S/N for dipole spatial correlations. The S/N of the dipole spatial correlation jumps from $\sim$1 to $\sim$3 when the 37th pulsar (PSR J2317+1439) is removed in the order of increasing noise.
@@ 5 Comparison with simulated data
To assess the peculiarity of the prominent features we see in our MCNMOS analysis of NG15, we make a comparison with an ensemble of simulated data sets that share the statistical properties of NG15. The simulated data were generated in Agazie et al. (2023a) following Pol et al. (2021). We note that the construction of these simulated data sets is central to the conclusions of this paper. Comprehensive details of their construction can be found in these previous references, but to summarize, each of the 100 simulated data sets share the same pulsars, timing cadence and noise parameters as NG15 and also include a spatially correlated power-law SGWB with power-law amplitude $2.7\times 10^{-15}$ and spectral index $-13/3$. The data sets differ only in their random-noise realizations. For each realization, we calculated the MCNMOS and removed pulsars one by one as in the analysis of NG15. By performing the same analysis on simulated data sets we are able to determine the expected behavior of the MCNMOS as pulsars are steadily removed. By comparing the results from the NG15 data with those from the ensemble of simulated data it is possible to assign a significance to the unexpected features identified in Section 4.3. Figure 6 shows the results for a single representative simulated data set with the least noisy pulsars removed first; this can be compared to the equivalent plot from the NG15 data set, shown in Figure 3.
@@ 5.1 Defining metrics
The metrics focus on S/N because each of our four features in Section 4.3 appear in S/N analyses. For completeness, however, we compute these metrics for $\hat{A}$ as well as the S/N. For each of these two variables, we calculate the three metrics for each of the three ORFs, resulting in a total of 18 metric values. These 18 values are calculated for NG15 and for the 100 simulations (see Figure 5, which shows the example of the distribution of the drop in S/N when the first pulsar is removed, for signals with HD spatial correlations). This allows us to compute $p$-values thereby determining the probability that a given NG15 pulsar removal feature would appear, given the null hypothesis that NG15 is accurately modeled and internally consistent. The 18 $p$-values are shown in Table 2; three of the 18 are less than or equal to 0.01. When calculating the metrics we only consider the data until the point where 50 pulsars have been removed; this is indicated in Figures 2, 3 and 6 by the gray regions. This cutoff is somewhat arbitrary, but is justified by the fact that after this point, the behavior of $\hat{A}^{2}$ and S/N becomes increasingly erratic because of the small number of low-sensitivity pulsars remaining. This is judged by eye from e.g., Figures 2, 3 and 4.
@@
Figure 4: The top panel shows the length of the timing baseline for each NG15 pulsar. The lower six panels are in pairs for each spatial correlation, tracking the MCNMOS and the S/N as the least noisy pulsars are removed one by one from the PTA. The panels with a red line correspond to monopole spatial correlations. The panels with a green line correspond to dipole spatial correlations. The panels with a blue line correspond to HD spatial correlations. The colored lines are the results from NG15 and each gray trace corresponds to one of the 100 simulated data sets. The black dashed lines show the medians of the 100 gray traces. The features described in Section 4.3 are highlighted.
@@ 6 Discussion
The behavior of the MCNMOS as applied to NG15, shown in Figure 3 and described in Section 4, exhibits several initially unexpected features as pulsars are steadily removed from the array. To contextualize the features and to determine whether they point to some unmodeled processes in the NG15, we have repeated the analysis using simulated data with known embedded GW signal and noise properties. To quantify the significance of the unexpected NG15 features we calculate $p$-values for 18 applications of the three metrics described in Section 5.1. Out of 18 instances, there are three with $p$-values less than or equal to 0.01.
@@
Dropout analysis (Aggarwal et al., 2019; Arzoumanian et al., 2020) was employed in NG15 to evaluate the significance of the GWB signal in individual pulsars. The dropout method introduces a binary parameter to either include or exclude the common signal for a specific pulsar. The dropout factor or the Bayes factor then indicates the degree to which each pulsar supports the presence of the GWB signal. In our approach, instead of evaluating each pulsar independently we removed pulsars in a cumulative manner. While both methods aim to understand the robustness and consistency of the detected signal, the dropout method focuses on the individual contribution of each pulsar in isolation. In contrast, our cumulative removal approach provides insight into the collective impact on the signal as pulsars are removed and so no single removal (other than the first) can be viewed in isolation. Comparing Figures 2 with 7 and 3 with 8 we see that the effect of removing any particular pulsar can be quite different depending on the order of removal. Furthermore, Figure 4 demonstrates that even when the order of removal is preserved, the removal of a particular pulsar can have a different effect on the recovered signal for different realizations of the simulation.
@@ 7 Conclusions and Future Work
We have investigated the behavior of recovered signal strength when pulsars are removed one by one from NG15. Despite variations during pulsar removal that we had not anticipated, these fluctuations are consistent with those observed in simulated data sets that have matching noise properties and a GWB injected. After calculating $p$-values for the data, we find that the results are consistent with the simulated data sets and, therefore, with an SGWB signal. The purpose of our analysis was to provide internal consistency checks on the NANOGrav 15 yr data set, and the results indicate that the data have successfully passed these tests. It also remains essential to conduct further investigations that may reveal systematics or unmodeled physical effects within the data.
The tools developed here have proved to be useful for checking the internal consistency of PTA data. Although we have applied them to NANOGrav 15 yr data set in this work, these tools could be applied routinely to new PTA data sets as they become available. We plan to analyze the EPTA DR2 data set next, with an eye on IPTA DR3 and MeerKAT data.
@@
An alphabetical order author list was used for this paper in recognition of the fact that a large, decade-timescale project such as NANOGrav is necessarily the result of the work of many people. All authors contributed to the activities of the NANOGrav Collaboration leading to the work presented here and reviewed the manuscript, text, and figures prior to the paper’s submission. Additional specific contributions to this paper are as follows. P.R.B., H.M., C.J.M., and A.V. collaboratively developed the concept and main ideas for this work. P.R.B. was responsible for data analysis and creating the figures. P.R.B., J.S.H., H.M., C.J.M., S.R.T., A.V., S.J.V., and C.A.W. contributed to writing the manuscript. N.S.P. created the simulated PTA data sets. G.A., A.A., A.M.A., Z.A., P.T.B., P.R.B., H.T.C., K.C., M.E.D., P.B.D., T.D., E.C.F., W.F., E.F., G.E.F., N.G., P.A.G., J.G., D.C.G., J.S.H., R.J.J., M.L.J., D.L.K., M. K., M.T.L., D.R.L., J.L., R.S.L., A.M., M.A.M., N.M., B.W.M., C.N., D.J.N., T.T.P., B.B.P.P., N.S.P., H.A.R., S.M.R., P.S.R., A.S., C.S., B.J.S., I.H.S., K.S., A.S., J.K.S., and H.M.W. developed the 15 yr data set through a combination of observations, arrival time calculations, data checks and refinements, and timing model development and analysis; additional specific contributions to the data set are summarized in Agazie et al. (2023b). J.S.H., J.G., N.L., N.S.P., J.P.S., and J.K.S. performed noise analyses on the data set.